Having difficulties with 3 problems Solving initial value problems with Laplace Transforms 1) ?′′−4?′−5?=3?^3, ?(0)=3, ?′(0)=3 2) ?′′−6? +9?=4−?(?−7), ?(0)=2, ?′(0)=0 and 3) Find the inverse Laplace Transform of (4s + 2)/(s^2 + 6s + 34)
Answer #1

1. The initial value problem is y -49-59= 3t3, y(0)=3,y/{09- 9 Taking Laplace toansform on both sides of the given differen3 1 3 56 e t 2 2 :. 419= 31{s:5}+{4}+i+Is I S+ 18 + I{34 84(8-5) (s+) (5-5) (5+) get test 84(8-5) (69 (1) By method of parti- 18 1) :.5B=D-4C 2-18 -AX 724 5 25 288 5 25 -18X5-288 25 - 378 25. 12) 2.B= 378 125 S3) 음대해 대학 3대해 1872 625 18722 25 125 18From equation (1), we have y(t) = zēt test 1872 3283 + 36 (2 - 2 3 4 625 19 est_zet 625 1872 625 125 25 378 +36 + 2 2 +² +628By method of partial fraction, we have sty - A + + BS (6-9) + cs. putting s =0,3, Wic Nawu A29A , AF A9-9 BS 42 3C cs 쌍m, A=AHence 9 where unit and u a heariside from equation (1), we have It) = 263t_bte3t+ 7 1 - e34-3test. I{EAS FC)} 19 test+9 - E